# The role of quantum and classical correlations in shrinking algorithms for optimization

### v1

@@ II-C Shrinking procedure
MaxCut is particularly suitable for shrinking algorithms, as there is a natural way of reducing the problem such that the shrunk problem is still a valid MaxCut problem. However, extensions of the shrinking presented here can be applied to other problems as well [12].
The shrinking procedure used in this work is similar to RQAOA [7, 11] as well as the algorithm introduced in Ref. [14]. For a better overview, a schematic of the algorithm is shown in Fig. 1. In general, the problem can be divided into three major steps a) to c). We will give a brief high-level overview before explaining all the steps in detail in the following paragraphs. After modeling the MaxCut problem, we first start with calculating correlations between the decision variables of the problem in step a). Then in step b), we use these correlations to successively reduce the number of nodes in the problem graph. We do this by combining two nodes, generating a new MaxCut instance that has one node fewer. The reduced problem is equivalent to the original problem under the additional constraint implied by the correlation. Importantly, since the shrunk problem is again a valid MaxCut problem, we can calculate new correlations for the shrunk problem using the same method as for the original problem. This can be done either after every shrinking step or after a given number $r$ of shrinking steps. Lastly, in step c), we recreate a solution to the original instance from a solution to the shrunk instance by undoing the shrinking steps. In the following, the three steps are described in more detail.
@@
At the core of our scheme is a method for computing the correlations between the decision variables of a given MaxCut problem instance. These are obtained by the routines presented in Section II-B. We want to emphasize that the shrinking procedure is independent of how the correlations are calculated. However, for different means of computing correlations, the algorithm’s performance will vary. This is because the correlations are calculated from approximations and will hence typically not be optimal. However, when using perfect correlations for every shrinking step, i.e., correlations obtained from an optimal cut of the problem, the shrinking algorithm is guaranteed to return the optimal cut.
@@
Next, we present how correlations are used to guide the shrinking procedure. After computing the correlations, we find the correlation with the largest absolute value, where ties are broken randomly. This way, the algorithm first selects the edges that have the strongest tendency to be cut or not cut.
Specifically, once the maximal correlation, say $b_{ij}$, is determined, the algorithm identifies both the vertex pair $i,j$ and the sign $\sigma_{ij}$ of the correlation:
@@
Then, for positive (negative) $\sigma_{ij}$, these two vertices are fixed to lie in the same (opposite) partition, reducing the total number of vertices by one (see panel b) in Fig. 1). To ensure consistency, the weights of the graph are updated such that the new graph $G^{\prime}=(V^{\prime},E^{\prime})$ is consistent with the original graph $(V,E)$ together with the newly imposed constraint for the nodes $i$ and $j$. For a node $i\in V$, its neighborhood is defined as $\mathcal{N}(i)\coloneqq\{k\in V\ |\ ik\in E\}$. For all neighbors $k\in\mathcal{N}(i)$ of node $i$, the new weights $\omega_{jk}^{\prime}$ are defined by
@@
while all other weights remain unaltered. In comparison to the original graph $(G,E)$, the updated graph $(G^{\prime},E^{\prime})$ lacks node $i$, and in some cases, two edges were merged into a single one:
@@
An example of these update rules is given in Fig. 1 b). It is important to keep track of the association of vertices in the shrunk graph with those of the original graph, in order to be able to reconstruct the final solution at the end.
On a more technical note, it is possible that after several shrinking steps, a node $u$ belonging to the considered correlation $b_{su}$ does not exist anymore because it had already been combined with another node $v$. This situation is depicted in panel b) of Fig. 1. There, we first calculate the correlations $b_{uv},b_{su},b_{ut},b_{tv}$ for the input graph on the left. The correlations are already sorted in descending order of absolute values. Then we shrink the graph on the left side of box b) according to the chosen correlation $b_{uv}$. This step yields the output graph on the right-hand side, where the node $u$ has been merged onto node $v$. The next best correlation is now $b_{su}$. However, node $u$ does not exist in the current graph. In this case, we simply interpret the correlation $b_{su}$ as the correlation $b_{sv}$. Note that we must take care of the sign used in the previous step to ensure consistency. To this end, we set $b_{sv}=\sigma_{uv}\cdot b_{su}$. Furthermore, analogous rules apply if both nodes $i$ and $j$ have already been combined with other nodes. For the special case that the two to-be shrunk nodes $i$ and $j$ already belong to the same node $k$, the shrinking step is skipped, and the algorithm continues by using the next best correlation. This skipping does not count as step in the procedure.
Since the shrinking step does not fix individual variables to be in a certain partition, but rather fixes the relation between two variables, the shrunk problem remains a valid MaxCut problem. This allows us to calculate correlations of the shrunk problem in the same manner as for the original graph. Intuitively, it is likely that correlations computed using the shrunk graph provide a better source of information for the shrinking step, as they have been specifically tailored to the shrunk graph. Therefore, it makes sense to introduce recalculations of the correlations for the shrunk graph after a given interval of $r$ shrinking steps, as also shown in Fig. 1 b). With these newly obtained correlations, we shrink the reduced problem for another $r$ steps until the correlations are recalculated again. For correlations obtained by QAOA and a recalculation interval of $r=1$, the shrinking algorithm corresponds to the RQAOA algorithm from Refs. [7, 11]. Note that recalculation can be performed using any of the means for computing the correlations discussed here.
In this work, the shrinking procedure is applied until the problem graph has been shrunk to a size of two nodes, when it becomes trivial to solve. A potential extension of our algorithm performs the shrinking procedure until the problem graph has been reduced to a sufficiently small size such that an exact classical solver can find an optimal solution.
@@
Finally, from a solution of the shrunk graph together with the history of all shrinking steps, a solution to the original graph is reconstructed. Starting from a solution of the shrunk graph, one simply has to backtrack through the performed shrinking steps. At each shrinking step, we either add the shrunk node to the vertex partition or exclude it, depending on the sign of the shrinking.
@@ III Results
Here, we present numerical experiments of applying the proposed shrinking algorithm using the different methods of computing correlations introduced in Sec. II. We first introduce the MaxCut  problem instances and performance metrics used in this work. Afterward, we compare the performance of the shrinking algorithm with the different means of calculating correlations to their corresponding bare approximation algorithms. Next, we investigate the influence of recalculating the correlations by studying the algorithm’s performance at different recalculation intervals. Finally, we employ tensor network methods to analyze the benefits of using deeper QAOA circuits to generate improved quantum correlations.
III-A Problem instances and approximation ratio for benchmarks
For the experiments in this study, two families of graphs are used. Most experiments are performed on Erdős-Rényi random graphs with a hundred nodes and different densities. An Erdős-Rényi graph for a given number of nodes and density $d$ is an undirected and unweighted graph with random edges. Each pair of nodes is connected by an edge with probability $d$.
Since simulating deep QAOA circuits requires computational resources that scale prohibitively with the graph density, we use sparse graphs for the tensor network experiments. Specifically, we use random 3-regular graphs with $50$ nodes. This type of graph has been thoroughly studied before in the context of QAOA for MaxCut  problems [39, 40, 41, 42] and has been used in hardware experiments because of their modest requirements on the qubit connectivity of quantum devices [43, 44].
Throughout the results section we use the approximation ratio as an indicator for the performance of the shrinking algorithm. We define the approximation ratio as
@@ III-A Problem instances and approximation ratio for benchmarks
Here, $S_{A}$ is the cut size obtained by the investigated algorithm. $S_{G}$ stands for the cut size retrieved from Gurobi [45], a state-of-the-art commercial solver. For each instance, we run Gurobi on a single core (Dual AMD Rome 7742) and terminate the optimization when the value of the optimization objective has not changed in an hour. We emphasize that in this way, it is not guaranteed that the Gurobi optimizer solves the problem instances to optimality, especially in the case of dense Erdős-Rényi graphs. However, we believe that the Gurobi benchmark is sufficiently strong for the purposes of this study.
@@ III-B Comparison to the bare underlying algorithm
We begin our analysis of the results by comparing the shrinking algorithm using various correlations to the approximate solutions obtained by the bare underlying algorithms used to compute the correlations. We first explain how the performance of the bare underlying algorithms is evaluated.
For the bare LP algorithm, we calculate the correlations for each edge in the manner described in Section II-B1 and then round these correlations to an integer solution by using a maximum spanning tree rounding heuristic. This is a well-established heuristic, previously used in Refs. [23, 46, 47]. For SDP, we use the standard Goemans-Williamson algorithm as the bare algorithm. Here, we do the hyperplane rounding (as described in Section II-B2) $15$ times and return the best solution. For QAOA, the graphs are too large to perform a classical simulation of the standard quantum algorithm. Efficient sampling from the output state of a QAOA $p=1$ circuit with one hundred qubits already exceeds the capabilities of classical hardware [48]. To this end, we calculate the expectation value of the energy returned by $p=1$ QAOA using analytical formulae from [30].
@@
Starting with LP, we observe that the shrinking with no recalculations outperforms the spanning tree heuristic for all of the considered graph densities, even though the correlations are identical in both cases. In addition, there is a clear improvement in performance when increasing the amount of recalculations performed (and, thus, decreasing $r$). This also holds true for SDP and QAOA. However, in contrast to SDP and QAOA, which show a roughly constant approximation ratio for all densities, LP performs significantly better at lower densities than on high densities.
@@
Finally, for QAOA the shrinking algorithm with recalculation interval $r=1$ (equivalent to RQAOA) sees a significant improvement over the bare algorithm: For the different densities, the median approximation ratio increases roughly from around $90\%$ for the bare QAOA to approximately $99\%$ for the $r=1$ shrinking algorithm. This behavior is in line with previous results from the literature [7, 11]. Interestingly, QAOA shows the largest relative improvement compared to the bare algorithm of the three different means of computing correlations. The bare algorithm and the shrinking with no recalculations perform quite similarly. However, the error bars indicate that the quality of the shrinking with no recalculations varies significantly more than the quality of the standard QAOA algorithm.
@@ III-C Effect of the recalculation interval and problem density
First, let us focus on the effect of the recalculation interval in general before looking at the impact of the problem density on the performance. For all three means of computing correlations, there is a clear trend that more recalculations improve the solutions. However, in the case of LP correlations, for low densities, the $r=10$ shrinking algorithm slightly outperforms its $r=1$ variant. While the performance differences for LP and the SDP between $r=10$ and $r=1$ are small, QAOA shows a significant increase for $r=1$. For all, SDP, LP and QAOA, the recalculation intervals $50$ and $\infty$ result in far worse approximation ratios than smaller recalculation intervals. This highlights the importance of recalculating correlations.
Turning now to the effect of problem density, the most striking dependence is observed for LP correlations. There, the approximation ratio for low-density instances is significantly better than for any other type of correlation. This is also in correspondence with the general good performance of the bare LP algorithm for sparse graphs, as also shown in other publications like Ref. [22].
Another notable feature in Fig. 3 is the systematic increase of the median approximation ratio for densities above $0.3$. This increase is likely an artifact of a worse quality of the reference solutions obtained via Gurobi, rather than improved performance of the shrinking algorithm. While for sparser instances, Gurobi can certify that the returned solution is optimal, the same is not true for denser instances. This leads us to believe that our choice of the benchmark is causing the increasing trend of the approximation ratio observed across all correlation types. However, it is important to note the present setup still allows us to perform comparisons between different correlation routines.
Thus, we see that the performance of the LP shrinking algorithm decreases very fast when the density increases from $0.05$. In contrast, the SDP and QAOA correlations lead to a more consistent performance of the shrinking algorithm across the entire range of densities. In addition, the SDP correlations clearly outperform $p=1$ QAOA for most densities, while the latter, in turn, outshines LP correlations for densities higher than $0.2$.
@@ III-D Performance of the GW correlations
After comparing SDP, LP and QAOA as means of computing correlations, we evaluate the GW correlations separately to ensure a fair comparison: While the previously analyzed strategies simply compute correlations using the problem graph, the algorithm for the GW correlations from paragraph II-B2 evaluates the MaxCut objective function several times directly while searching for a good hyperplane. Thus, it utilizes information not accessible to the other routines for computing correlations. Nonetheless, these correlations can still be computed efficiently in polynomial time.
@@
Furthermore, the GW correlations for $r=50$ have a similar performance as the SDP correlations for $r=1$. From this, we conclude that the best approximation ratio for the SDP correlations can also be achieved by the GW correlations with less computational resources. Finally, it is worth mentioning that GW correlations with a recalculation interval of $r=10$ yield a better median approximation ratio than with $r=1$ for all densities. The reason for this is that the GW correlations use additional information, which has a larger positive impact on the performance when correlations are used for several shrinking steps without recalculations. This finding is particularly interesting since less computational resources yield a better result, in contrast to the analysis for the SDP correlations in the previous chapter. There, the performance steadily increases with the number of recalculations. As discussed more in detail in Section IV, this behaviour could result from the additional information about the cut value to which only the GW correlations have access.

### v2

@@ II-C Shrinking procedure
MaxCut is well-suited for shrinking algorithms since there is a natural way of reducing the problem such that the shrunk problem is still a valid MaxCut problem. However, the shrinking procedure presented here can be extended to other problems as well [12].
The shrinking procedure used in this work is similar to RQAOA [7, 11] and the algorithm introduced in Ref. [14]. Fig. 1 presents a schematic of the proposed Algorithm 1.
Algorithm 1 Shrinking algorithm
Input: Graph $G=(V,E)$,
                Recalculation interval $r$,
                Steps $N$, $\triangleright$ Number of shrinking steps
      Output: Solution $S$.
1:G’ $\leftarrow$ G
[…]
3:
4:for step in $\{0,...,N-1\}$ do
5:     if step mod r is 0 then $\triangleright$ Recalculate every r steps
6:         Cor $\leftarrow$ Calculate correlations for G’.
                   See section II-B and Fig. 1 (a).
7:         Cor $\leftarrow$ Sort correlations by absolute value.
8:     end if
9:     $(i,j)$, $\sigma_{ij}$leftarrow Pop first element of Cor.
10:     if i or j already collapsed onto other node then
11:         $(i,j)$, $\sigma_{ij}\leftarrow$ Ensure consistency with G’.
                          Refer to text for further details.
12:     end if
13:     G’ $\leftarrow$ shrink edge(G’, (i,j), $\sigma_{ij}$) $\triangleright$ See algorithm 2
14:                                                   and Fig. 1 (b)
15:     Append shrinking step $\{(i,j),\sigma_{ij}\}$ to L.
16:end for
17:
18:S’ $\leftarrow$ Solve MaxCut on shrunk graph G’.
19:S $\leftarrow$ Reconstruct solution from solution $S^{\prime}$ and history L.           $\triangleright$ See Fig 1 (c)
20:return $S$triangleright Return Solution
It can be divided into three major steps. We will give a brief overview before explaining the individual steps in detail. For a given MaxCut problem, we first calculate correlations between the decision variables (step (a) in Fig. 1). Then, we use these correlations to successively reduce the number of nodes in the problem graph (step (b) in Fig. 1). To this end, we combine two nodes, generating a new MaxCut instance where the number of nodes is decreased by one. The reduced problem is equivalent to the original problem with the additional constraint implied by the correlation. Importantly, since the shrunk problem is again a MaxCut problem, we can calculate new correlations for the shrunk problem using the same method as for the original problem. We re-calculate correlations after a given number $r$ of shrinking steps. Finally, we recreate a solution to the original instance from a solution to the shrunk instance by undoing the shrinking (step (c) in Fig. 1). In the following, we describe the three steps in detail.
@@
The core of our algorithm is a method for computing correlations between decision variables of a given MaxCut instance. We employ the methods presented in Section II-B. However, the shrinking algorithm 1 can be used with any method for calculating correlations. Importantly, its performance depends on the quality of the employed correlations, defined as the value of the best cut fulfilling these correlations. In particular, the shrinking algorithm is guaranteed to return an optimal cut if the correlations are obtained from an optimal cut.
@@
Next, we present how correlations are used to guide the shrinking procedure. A pseudocode representation of the following explanation can be found in algorithms 1 and 2.
Algorithm 2 Shrink edge
Input: Graph $G$,
                Correlation - edge $(i,j)$, sign $\sigma_{ij}$,
      Output: Graph G’ with one less node than G.
1:E’ $\leftarrow$ get new edges from eq. (18).
2:Initialize empty weights $\omega^{\prime}$.
3:for $(j,k)$ in E’ do $\triangleright$ Update edge weights
4:     if (j, k) in E then
5:         $\omega^{\prime}_{jk}\leftarrow\omega_{jk}+\sigma_{ij}\omega_{ik}$
6:     else
7:         $\omega^{\prime}_{jk}\leftarrow\sigma_{ij}\omega_{ik}$
8:     end if
9:end for
10:G’ $\leftarrow$ Create graph from $\omega^{\prime}$.
11:return G’ $\triangleright$ Return graph with one less node
After computing correlations, we choose the correlation with the largest absolute value, say $b_{ij}$, where ties are broken randomly. The rationale is that we select edges first that have a strong tendency to be cut or not cut. Then, we consider the sign $\sigma_{ij}$ of the correlation,
@@
For positive (negative) $\sigma_{ij}$, the vertices $i$ and $j$ are fixed to lie in the same (opposite) partition, reducing the total number of vertices by one, see panel b) in Fig. 1. To ensure consistency, we update the weights of new graph $G^{\prime}=(V^{\prime},E^{\prime})$ in the following way. For a node $i\in V$, we define its neighborhood as $\mathcal{N}(i)\coloneqq\{k\in V\ |\ ik\in E\}$. For all nodes $k\in\mathcal{N}(i)$, we define new weights $\omega_{jk}^{\prime}$ by
@@
while all other weights remain unchanged. In comparison to the original graph $(G,E)$, the updated graph $(G^{\prime},E^{\prime})$ lacks node $i$ and some edges were merged,
@@
An example of these update rules is given in Fig. 1 b) and the pseudocode for generating this updated graph can be found in Algorithm 2. It is important to keep track of the imposed correlations in order to reconstruct the final solution at the end.
On a more technical note, it is possible that after several shrinking steps, a node $u$ belonging to the considered correlation $b_{su}$ does not exist anymore because it had already been combined with another node $v$. This situation is shown in panel b) of Fig. 1. In this case, we simply interpret the correlation $b_{su}$ as the correlation $b_{sv}$. Note however, that we must take care of the sign used in the previous step to ensure consistency. To this end, we set $b_{sv}=\sigma_{uv}\cdot b_{su}$. Furthermore, analogous rules apply if both nodes $i$ and $j$ have already been combined with other nodes. For the special case that the two nodes $i$ and $j$ to be shrunk already belong to the same node $k$, the shrinking step is skipped, and the algorithm continues by using the next best correlation. This skipping does not count as step in the procedure.
Since the shrinking step does not fix individual variables to be in a certain partition, but rather fixes the relation between two variables, the shrunk problem remains a valid MaxCut problem. This allows us to calculate correlations of the shrunk problem in the same manner as for the original graph. Intuitively, correlations computed for the shrunk graph provide a better source of information for the next shrinking step than re-using correlations for the original graph. Therefore, it makes sense to introduce recalculations of the correlations for the shrunk graph after a given interval of $r$ shrinking steps, as shown in Fig. 1 b) and Algorithm 1. With these newly obtained correlations, we shrink the reduced problem for another $r$ steps. For correlations obtained by QAOA and a recalculation interval of $r=1$, the shrinking algorithm corresponds to the RQAOA algorithm from Refs. [7, 11].
In this work, we apply the shrinking procedure until the problem graph has only two nodes and is trivial to solve. A potential extension of our algorithm performs the shrinking procedure until the problem graph has been reduced to a sufficiently small size such that an exact classical solver can find an optimal solution.
@@
Finally, from a solution of the shrunk graph together with the history of all shrinking steps, we reconstruct a solution to the original graph, see Algorithm 1. Starting from a solution of the shrunk graph, we simply undo the performed shrinking steps.
@@ III Results
Here, we present experimental results for the proposed shrinking algorithm using different methods for computing correlations, introduced in Sec. II-B. We first introduce the considered MaxCut instances and performance metrics. Then, we compare the performance of the shrinking algorithm with the different means of calculating correlations to the corresponding bare approximation algorithms. Next, we investigate the influence of recalculating the correlations by studying the algorithm’s performance at different recalculation intervals. Finally, we employ tensor network methods to analyze the benefits of using deeper QAOA circuits to generate improved quantum correlations.
III-A Problem instances and performance metric
We consider two families of graphs. In most of our experiments, we solve MaxCut on Erdős-Rényi random graphs with one hundred nodes and different densities. For a given number of nodes and density $d$, an Erdős-Rényi graph is an undirected and unweighted graph where each pair of nodes is connected by an edge with probability $d$.
Since simulating deep QAOA circuits requires computational resources that scale prohibitively with the graph density, we use sparse graphs for the tensor network experiments. Specifically, we use random 3-regular graphs with $50$ nodes. This graph ensemble has been studied before in the context of QAOA for MaxCut  problems [39, 42, 43, 44] and has been used in hardware experiments because of the modest requirements on the qubit connectivity of quantum devices [45, 46].
Throughout the results section we use the approximation ratio as the performance metric. We define the approximation ratio as
@@ III-A Problem instances and performance metric
Here, $S_{A}$ is the cut size obtained by the investigated algorithm and $S_{G}$ is the cut size retrieved from Gurobi [47], a state-of-the-art solver. For each instance, we run Gurobi on a single core (Dual AMD Rome 7742) and terminate the optimization when the objective value has not changed in an hour. We emphasize that in this way, it is not guaranteed that Gurobi solves the problem instances to optimality. However, we believe that the Gurobi benchmark is sufficiently strong for the purposes of this study.
@@ III-B Comparison to the bare underlying algorithm
We begin our analysis of the results by comparing the shrinking algorithm using various correlations to the solutions obtained by the bare underlying algorithms used to compute the correlations. We first explain how the performance of the bare underlying algorithms is evaluated.
For the bare LP algorithm, we calculate the correlation for each edge in the manner described in Section II-B1. Then, we round these correlations to an integer solution by using a rounding heuristic based on a maximum-weight spanning-tree computation. This is a well-established heuristic, previously used in Refs. [21, 48, 49]. For SDP, we use the standard Goemans-Williamson algorithm as the bare algorithm. Here, we perform the hyperplane rounding (as described in Section II-B2) $15$ times and return the best solution. For QAOA, the graphs are too large to perform a classical simulation of the quantum algorithm. Efficient sampling from the output state of a QAOA $p=1$ circuit with one hundred qubits already exceeds the capabilities of classical hardware [50]. Therefore, we calculate the expectation value of the energy returned by $p=1$ QAOA using the analytical formulae from [30].
@@
Starting with LP, we observe that the shrinking with no recalculations outperforms the spanning tree heuristic for all of the considered graph densities, even though the correlations are identical in both cases. In addition, there is a clear improvement in performance when increasing the number of recalculations, that is, when decreasing $r$. The same holds true for SDP and QAOA. However, LP performs significantly better at lower densities than at high densities while SDP and QAOA show a roughly constant approximation ratio for all densities.
@@
Finally, for QAOA the shrinking algorithm with recalculation interval $r=1$ (equivalent to RQAOA) sees a significant improvement over the bare algorithm. For all densities, the median approximation ratio increases roughly from around $90\%$ for the bare QAOA to approximately $99\%$ for the $r=1$ shrinking algorithm. This behavior is in line with previous results from the literature [7, 11]. Interestingly, QAOA shows the largest relative improvement compared to the bare algorithm among the three different means of computing correlations. The bare algorithm and the shrinking with no recalculations perform quite similarly. However, the error bars indicate that the quality of the shrinking with no recalculations varies significantly more than the quality of the standard QAOA algorithm.
@@ III-C Effect of the recalculation interval and problem density
First, we focus on the effect of the recalculation interval before studying the impact of the problem density on the performance. For all three means of computing correlations, we observe a clear trend that more recalculations improve the solutions. However, in the case of LP correlations, for low densities, the $r=10$ shrinking algorithm slightly outperforms the $r=1$ variant. While for LP and the SDP the performance differences between $r=10$ and $r=1$ are small, QAOA shows a significant increase from $r=10$ to $r=1$. For SDP, LP and QAOA, the recalculation intervals $r=50$ and $r=\infty$ result in far worse approximation ratios than smaller recalculation intervals. This highlights the importance of recalculating correlations.
Turning to the effect of problem density, the most striking dependence is observed for LP correlations. There, the approximation ratio for low-density instances is significantly better than for any other type of correlation. This is in agreement with the good performance of LP-based branch-and-bound algorithms for sparse MaxCut instances [20, 24].
A general feature in Figs. 3 a) - c) is the systematic increase of the median approximation ratio for densities above $0.3$. This increase is likely an artifact caused by a decreasing quality of the reference solutions obtained via Gurobi, rather than an improved performance of the shrinking algorithm. While for sparse instances, Gurobi can certify that the returned solution is optimal, the same is not true for denser instances. Thus, we believe that our choice of the benchmark metric causes the increasing trend in the approximation ratio observed across all correlation types. However, our benchmark still allows us to compare between different correlation routines.
Thus, we observe that the performance of the LP shrinking algorithm decreases rapidly when the density increases from $0.05$. In contrast, the SDP and QAOA correlations show to a more consistent performance of the shrinking algorithm across the entire range of densities. Moreover, the SDP correlations clearly outperform $p=1$ QAOA for most densities, while the latter, in turn, outperforms LP correlations for densities above $0.2$.
@@ III-D Performance of the GW correlations
After comparing SDP, LP and QAOA as means of computing correlations, we evaluate the GW correlations separately to ensure a fair comparison: While the previously analyzed strategies simply compute correlations using the problem graph, the algorithm for the GW correlations from paragraph II-B2 evaluates the MaxCut objective function several times directly while searching for a good hyperplane. Thus, it utilizes information not accessible to the other routines for computing correlations. Nonetheless, these correlations can still be computed efficiently.
@@
Furthermore, the GW correlations for $r=50$ have a similar performance as the SDP correlations for $r=1$. From this, we conclude that the best approximation ratio for the SDP correlations can also be achieved by the GW correlations with less computational resources. Finally, it is worth mentioning that GW correlations with a recalculation interval of $r=10$ yield a better median approximation ratio than with $r=1$ for all densities. It seems that the additional information about a high-quality MaxCut solution, e.g. the well chosen hyperplane, is most beneficial when correlations are used for several shrinking steps without recalculations. Refer to section IV for a further discussion of this observation. This finding is particularly interesting since less computational resources yield a better result, in contrast to the analysis for the SDP correlations in the previous chapter. There, the performance steadily increases with the number of recalculations. As discussed more in detail in Section IV, this behavior could result from the additional information about the cut value to which only the GW correlations have access.